# RM-R1: Reward Modeling as Reasoning

### v2

@@ Abstract
Reward modeling is essential for aligning large language models (LLMs) with human preferences through reinforcement learning (RL). To provide accurate reward signals, a reward model (RM) should stimulate deep thinking and conduct interpretable reasoning before assigning a score or a judgment. Inspired by recent advances of long chain-of-thought (CoT) on reasoning-intensive tasks, we hypothesize and validate that integrating reasoning capabilities into reward modeling significantly enhances RM’s interpretability and performance. To this end, we introduce a new class of generative reward models – Reasoning Reward Models (ReasRMs) – which formulate reward modeling as a reasoning task. We propose a reasoning-oriented training pipeline and train a family of ReasRMs, RM-R1. RM-R1 features a chain-of-rubrics (CoR) mechanism – self-generating sample-level chat rubrics or math/code solutions, and evaluating candidate responses against them. The training of RM-R1 consists of two key stages: (1) distillation of high-quality reasoning chains and (2) reinforcement learning with verifiable rewards. Empirically, our models achieve state-of-the-art performance across three reward model benchmarks on average, outperforming much larger open-weight models (e.g., INF-ORM-Llama3.1-70B) and proprietary ones (e.g., GPT-4o) by up to 4.9%. Beyond final performance, we perform thorough empirical analysis to understand the key ingredients of successful ReasRM training. To facilitate future research, we release six ReasRM models along with code and data at https://github.com/RM-R1-UIUC/RM-R1.
@@ 1 Introduction
Reward models (RMs) play a critical role in LLM post-training, particularly in reinforcement learning with human feedback (RLHF) [4, 23], where they serve as scalable proxies for human evaluators. Existing research on reward modeling can be broadly classified into two categories: (1) scalar-based reward models (ScalarRM)  [19] and (2) generative reward models (GenRM)  [37]. Scalar-based approaches frame reward modeling as a classification problem, typically training a sequence classifier on top of a language model. In contrast, generative approaches retain the original language model decoding head and leverage the model’s generative abilities to produce free-form pairwise judgments. While scalar-based methods are direct and often effective, they are opaque, offering no intermediate reasoning steps to justify the model’s decisions. This lack of transparency may limit their capacity to handle more challenging, reasoning-intensive preference tasks. On the other hand, although generative methods provide greater transparency, their reasoning is often superficial and unhelpful for reliable judgment, leading to suboptimal performance [7, 21].
@@
In this work, we unleash the reasoning potential of RMs and propose a new class of models: Reasoning Reward Models (ReasRMs). Different from standard GenRMs, ReasRMs emphasize leveraging long and coherent reasoning chains during the judging process to enhance the model’s ability to assess and distinguish complex outputs accurately. We hypothesize and validate that integrating long reasoning chains during the judging process significantly enhances downstream reward model performance. We explore several strategies for adapting instruction-tuned language models into logically coherent ReasRMs. Notably, we find that solely applying reinforcement learning with verifiable rewards (RLVR) [12] in reward modeling does not fully realize the model’s reasoning capabilities. We also observe that plain chain-of-thought (CoT) reasoning falls short at perceiving the fine-grained distinction across different question types.
Through a series of studies, we design a training pipeline that introduces reasoning distillation prior to RLVR, ultimately resulting in the development of RM-R1. To guide RM-R1 for structured rollouts tailored for reward modeling during RL training and inference, we design a Chain-of-Rubrics (CoR) process. Specifically, the model categorizes the input sample into one of two categories: chat or reasoning. For chat tasks, the model generates a set of evaluation rubrics, justifications for the rubrics, and evaluations tailored to the specific question. For reasoning tasks, correctness is the most important and generally preferred rubrics, so we directly let the model first solve the problem itself before evaluating and picking the preferred response. This task perception enables RLVR to fully exploit reasoning supervision for reasoning-heavy tasks while producing coherent, high-quality evaluation chains for chat tasks. In addition, we explore how to directly adapt existing reasoning models into reward models. Since these models have already undergone substantial reasoning-focused distillation, we fine-tune them using RLVR without additional distillation stages. Based on our training recipes, we produce RM-R1 models ranging from 7B to 32B.
Empirically, RM-R1 models consistently yield highly interpretable and coherent reasoning traces. On average, RM-R1 achieves state-of-the-art performance on RewardBench [16], RM-Bench [20], and RMB [39], outperforming 70B, 340B, GPT-4o, and Claude models by up to 4.9% in accuracy. Beyond final performance, we conduct extensive empirical analyses of RM-R1, including studies of its scaling effects, ablations of our training recipes, comparisons with non-reasoning baselines, training dynamics, and detailed case studies. In summary, our main contributions are as follows:
@@ 2 RM-R1
Figure 2 presents the overall training pipeline of RM-R1, which consists of two stages: reasoning distillation and reinforcement learning. (1) Reasoning Distillation: Starting from an off-the-shelf instruction-tuned model (e.g., Qwen-2.5-14B-Instruct), we further train the model using synthesized high-quality reasoning traces. This stage equips RM-R1 with essential reasoning capabilities required for effective reward modeling. (2) Reinforcement learning: While distillation is effective for injecting reasoning patterns, distilled models often overfit to specific patterns in the training data, limiting their generalization ability. To overcome this limitation, we introduce a reinforcement learning phase that further optimizes the model, resulting in the final version of RM-R1.
@@ 2.1 Task Definition
$r_{\theta}(j|x,y_{a},y_{b})=\prod_{t=1}^{T}r_{\theta}(j_{t}|x,y_{a},y_{b},j_{<t}).$
@@ 2.2 Reasoning Distillation for Reward Modeling
For an instruction-tuned model (e.g., Qwen-2.5-14b-instruct [35]), it is quite intuitive to turn it into a GenRM simply by prompting. However, without fine-tuning on reward modeling reasoning traces, these models may struggle to conduct consistent judgments. To bootstrap its reasoning potential, we start with training an instruction-tuned model with long reasoning traces synthesized for reward modeling. Specifically, we sample $M$ data samples from $\mathcal{D}$ and denote it as $\mathcal{D}_{\rm sub}$. Given a data sample $(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\in\mathcal{D}_{\rm sub}$, we ask an “oracle” model like o3 or claude-3-7-sonnet to generate its structured reasoning trace $r^{(i)}$ justifying why $y_{l}^{(i)}$ is chosen as the correct answer of $x^{(i)}$. We then construct the reasoning trace ground truth:
@@
Formally, the objective of distillation is to adjust $\theta$ to maximize the likelihood of generating the desired reasoning trace and response $y$ given the prompt $x$. We minimize the negative log-likelihood (NLL) loss:
$\mathcal{L}_{\mathrm{distill}}(\theta)=-\sum_{(x,y)\in\mathcal{D}_{\mathrm{distill}}}\sum_{t=1}^{T}\log r_{\theta}\left(y_{t}\mid x,y_{<t}\right),$ (5)
@@ 2.3 RL Training
where $r_{\mathrm{ref}}$ is the reference reward model. In practice, we use the checkpoint before training as $r_{\mathrm{ref}}$, and that means $r_{\mathrm{ref}}$ could be an off-the-shelf LLM or the LLM obtained after the distillation step in Section 2.2. $\mathcal{R}(x,j)$ is the reward function, and $\mathbb{D}_{\mathrm{KL}}$ is KL-divergence. The $x$ denotes input prompts drawn from the preference data $\mathcal{D}$. The $j$ indicates the text generated by the reward model, which includes the reasoning trace. In practice, we use Group Relative Policy Optimization (GRPO) [27] to optimize the objective in Equation 6, the details of which can be find in Appendix C.
@@ 2.3.1 Chain-of-Rubrics (CoR) Rollout
To facilitate the distilled models to proactively generate critical reasoning traces, we design a system prompt as shown in Figure 3 during rollout. Intuitively, reward modeling for general domain (e.g., chat, safety, etc.) and reasoning domain (e.g., math, code, etc.) should focus on different angles. For example, for the chat domain, we may care more about some aspects that can be expressed in textual rubrics (e.g., be polite), yet for the reasoning domain, we usually care more about logical coherence and answer correctness. Based on this intuition, we instruct ${r_{\theta}}$ to classify each preference data sample $\{(x,y_{c},y_{r})\}$ into one of the two <type>: Chat or Reasoning. For each <type>, we prompt ${r_{\theta}}$ to carry out the behavior corresponding to that type step by step: For reasoning tasks, we ask ${r_{\theta}}$ to solve $x$ on its own. During the <eval> phase, $r_{\theta}$ compares $y_{c}$ and $y_{r}$ conditioned on its own </solution> and selects an <answer>. Regarding the Chat type, we instead ask ${r_{\theta}}$ to think about and justify the <rubric> for grading the chat quality (including safety). Since Chat scenarios do not have a verifiable reward, we instruct the model to select the <answer> based on its own <rubric>.
@@ 3.2 Main Results
State-of-the-Art Performance. Our RM-R1-DeepSeek-Distilled-Qwen-14B model surpasses all previous leading Reward Models (RMs), including INF-ORM-Llama3.1-70B, Nemotron-4-340B-Reward, and GPT-4o, while operating at a much smaller scale on average. Our 32B models, RM-R1-Qwen-Instruct-32B and RM-R1-DeepSeek-Distilled-Qwen-32B, further extend this performance advantage by a non-trivial margin. The success of RM-R1 is attributable to both our meticulously designed training methodology and the effective scaling of our models, as extensively analyzed in Section 4.1 and Section 4.2. In particular, RM-R1 outperforms existing top-tier ScalarRMs. This highlights the considerable potential of ReasRMs, a category where prior GenRMs have exhibited suboptimal performance and are generally not comparable to their scalar counterparts. Simultaneously, our comprehensive evaluation indicates that the top-performing scalar models on RewardBench do not consistently achieve state-of-the-art (SOTA) performance; in fact, larger models frequently underperform smaller ones. This evaluation underscores the need for a more comprehensive and systematic approach to RM assessment.
@@ 4.2 Scaling Effects
We then investigate how model performance varies with scale, considering both model size and inference-time compute. In some cases – such as scalar-valued sequence models from Skywork [19] and InternLM2 [6] – the smaller models (8b/7b) outperforms the larger ones (27b/20b), showing no advantage of scaling. In this subsection, we show that this trend does not hold for RM-R1, where scaling brings clear and substantial improvements.
@@ 4.2.2 Inference-time Computation
To ensure a fair comparison, we match the training rollout budget to the inference budget in each setting (i.e., we allow a maximum of $x$ tokens during training for a compute budget of $x$ at inference). All models are trained using GRPO with identical datasets and hyperparameter configurations. Figure 4(b) shows the relationship between compute budget and performance. We observe a clear improvement trend as the inference budget increases. This highlights the benefits of long reasoning chains in reward modeling.
@@ 4.3 Effectiveness of Reasoning Training
Instruct + Distilled + SFT. This approach applies SFT (with respect to the answer directly) after the distillation stage, serving as a direct comparison point with RM-R1.
Instruct + RM-R1 (Distilled + RL). This is the full approach proposed in this paper, following the procedure detailed in Section 2 and Section 4.1.
@@ 4.4 Case Study
Table 4: Case Study of RM-R1(succinct version). Complete model responses are shown in Table 9 and Table 10. Cold-Start RL evaluation tends to focus on superficial features (e.g., number of suggestions, relevance), whereas RM-R1 produces high-quality rubrics (e.g., accuracy for factual questions), faithfully follows them, and evaluates responses based on their actual content.
@@ 6 Conclusion and Future Work
The research reported in this paper/presentation was performed in connection with the U.S. Army Contracting Command - Aberdeen Proving Ground (ACC-APG) and the Defense Advanced Research Projects Agency (DARPA) under contract number W912CG- 24-C-0011. The views and conclusions in this paper/presentation are those of the authors and should not be interpreted as presenting the official policies or position, either expressed or implied, of ACC-APG, DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation hereon.
@@ D.3 Baselines
Scalar RMs. Scalar RMs produce a score for model response directly, predicting preference through single numeric scores without explicit reasoning traces. This category includes models such as RM [30], SteerLM-RM [34], Nemotron-RM [2], Starling-RM [40], ArmoRM [32], Skywork-RM [19], etc. These models typically show strong performance in clearly defined tasks but often lack interpretability and nuanced reasoning capability.
GenRMs. Generative reward models provide more nuanced feedback by generating textual judgments. Examples include Llama [11], Qwen [35], Claude [3], GPT-4 variants [1, 14], Gemini-1.5-pro [24], and Self-taught Evaluators [33]. GenRMs leverage generative capabilities of LLMs to produce explanations and rationales alongside numerical scores, enhancing interpretability and decision transparency.
ReasRMs. Reasoning-enhanced reward models explicitly incorporate reasoning processes into their judgments, often trained through critiques or chain-of-thought strategies. Notable examples are Critique-RM [36], DeepSeek-GRM [21], and our proposed RM-R1 models. These models excel in tasks demanding rigorous reasoning, safety evaluations, and nuanced preference judgments due to their grounding in structured critical thinking.

### v3

@@ Abstract
Reward modeling is essential for aligning large language models with human preferences through reinforcement learning from human feedback. To provide accurate reward signals, a reward model (RM) should stimulate deep thinking and conduct interpretable reasoning before assigning a score or a judgment. Inspired by recent advances of long chain-of-thought on reasoning-intensive tasks, we hypothesize and validate that integrating reasoning capabilities into reward modeling significantly enhances RM’s interpretability and performance. To this end, we introduce a new class of generative reward models – Reasoning Reward Models (ReasRMs) – which formulate reward modeling as a reasoning task. We propose a reasoning-oriented training pipeline and train a family of ReasRMs, RM-R1. RM-R1 features a chain-of-rubrics (CoR) mechanism – self-generating sample-level chat rubrics or math/code solutions, and evaluating candidate responses against them. The training of RM-R1 consists of two key stages: (1) distillation of high-quality reasoning chains and (2) reinforcement learning with verifiable rewards. Empirically, our models achieve state-of-the-art performance across three reward model benchmarks on average, outperforming much larger open-weight models (e.g., INF-ORM-Llama3.1-70B) and proprietary ones (e.g., GPT-4o) by up to 4.9%. Beyond final performance, we perform thorough empirical analyses to understand the key ingredients of successful ReasRM training. To facilitate future research, we release six ReasRM models along with code and data at https://github.com/RM-R1-UIUC/RM-R1.
@@ 1 Introduction
Reward models (RMs) play a critical role in large language model (LLM) post-training, particularly in reinforcement learning with human feedback (RLHF) [4, 24], where they serve as scalable proxies for human evaluators. Existing research on reward modeling can be broadly classified into two categories: (1) scalar-based reward models (ScalarRM)  [20] and (2) generative reward models (GenRM)  [40]. Scalar-based approaches frame reward modeling as a classification problem, typically training a sequence classifier on top of a language model. In contrast, generative approaches retain the original language model decoding head and leverage the model’s generative abilities to produce free-form pairwise judgments. While scalar-based methods are direct and often effective, they are opaque, offering no intermediate reasoning steps to justify the model’s decisions. This lack of transparency may limit their capacity to handle more challenging, reasoning-intensive preference tasks. On the other hand, although generative methods provide greater transparency, their reasoning is often superficial and unhelpful for reliable judgment, leading to suboptimal performance [7, 22].
@@
In this work, we unleash the reasoning potential of RMs and propose a new class of models: Reasoning Reward Models (ReasRMs). Different from standard GenRMs, ReasRMs emphasize leveraging long and coherent reasoning chains during the judging process to enhance the model’s ability to assess and distinguish complex outputs accurately. We validate that integrating long reasoning chains during the judging process significantly enhances downstream reward model performance. We explore several strategies for adapting instruction-tuned language models into logically coherent ReasRMs. Notably, we find that solely applying reinforcement learning with verifiable rewards (RLVR) [12] in reward modeling does not fully realize the model’s reasoning capabilities. We also observe that plain chain-of-thought (CoT) reasoning falls short at perceiving the fine-grained distinction across different question types.
Through a series of studies, we design a training pipeline that introduces reasoning distillation prior to RLVR, ultimately resulting in the development of RM-R1. To fully elicit the reasoning capability of RM-R1 for reward modeling, we design a Chain-of-Rubrics (CoR) process. Specifically, the model categorizes the input sample into one of two categories: chat or reasoning. For chat tasks, the model generates a set of evaluation rubrics, justifications for the rubrics, and evaluations tailored to the specific question. For reasoning tasks, correctness is the most important and generally preferred rubrics, so we directly let the model first solve the problem itself before evaluating and picking the preferred response. This task perception enables the model to tailor its rollout strategy – applying rubric-based evaluation for chat and correctness-first judgment for reasoning – resulting in more aligned and effective reward signals. In addition, we explore how to directly adapt existing reasoning models into reward models. Since these models have already undergone substantial reasoning-focused distillation, we fine-tune them using RLVR without additional distillation stages. Based on our training recipes, we produce RM-R1 models ranging from 7B to 32B.
Empirically, RM-R1 models consistently yield highly interpretable and coherent reasoning traces. On average, RM-R1 achieves state-of-the-art performance on RewardBench [17], RM-Bench [21], and RMB [43], outperforming 70B, 340B, GPT-4o, and Claude models by up to 4.9%. Beyond final performance, we conduct extensive empirical analyses of RM-R1, including ablations of our training recipes, studies of its scaling effects, comparisons with non-reasoning baselines, detailed case studies, and training dynamics. In summary, our main contributions are as follows:
@@ 2 RM-R1
Figure 2 presents the overall training pipeline of RM-R1, which consists of two stages: reasoning distillation and reinforcement learning. (1) Reasoning Distillation: Starting from an off-the-shelf instruction-tuned model (e.g., Qwen-2.5-14B-Instruct), we further train the model using synthesized high-quality reasoning traces. This stage equips RM-R1 with essential reasoning capabilities required for effective reward modeling. (2) Reinforcement learning: While distillation is effective for injecting reasoning patterns, distilled models often overfit to specific patterns in the training data, limiting their generalization ability [9]. To overcome this limitation, we introduce a reinforcement learning phase that further optimizes the model, resulting in the final version of RM-R1.
@@ 2.1 Task Definition
$r_{\theta}(j|x,y_{a},y_{b})=\prod_{t=1}^{T}r_{\theta}(j_{t}|x,y_{a},y_{b},j_{<t}).$ (2)
@@ 2.2 Reasoning Distillation for Reward Modeling
For an instruction-tuned model (e.g., Qwen-2.5-14b-instruct [37]), it is quite intuitive to turn it into a GenRM simply by prompting. However, without fine-tuning on reward modeling reasoning traces, these models may struggle to conduct consistent judgments. To bootstrap its reasoning potential, we start with training an instruction-tuned model with long reasoning traces synthesized for reward modeling. Specifically, we sample $M$ data samples from $\mathcal{D}$ and denote it as $\mathcal{D}_{\rm sub}$. Given a data sample $(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\in\mathcal{D}_{\rm sub}$, we ask an “oracle” model like o3 or claude-3-7-sonnet to generate its structured reasoning trace $r^{(i)}$ justifying why $y_{l}^{(i)}$ is chosen as the preferred response of $x^{(i)}$. We then construct the reasoning trace ground truth:
@@
Formally, the objective of distillation is to adjust $\theta$ to maximize the likelihood of generating the desired reasoning trace and picking the response $y$ given the prompt $x$. We minimize the negative log-likelihood (NLL) loss:
$\mathcal{L}_{\mathrm{distill}}(\theta)=-\sum_{(x,y)\in\mathcal{D}_{\mathrm{distill}}}\sum_{t\in[|y|]}\log r_{\theta}\left(y_{t}\mid x,y_{<t}\right),$ (6)
@@ 2.3 RL Training
where $r_{\mathrm{ref}}$ is the reference reward model. In practice, we use the checkpoint before RL training as $r_{\mathrm{ref}}$, and that means $r_{\mathrm{ref}}$ could be an off-the-shelf LLM or the LLM obtained after the distillation step in Section 2.2. $\mathcal{R}(x,j)$ is the reward function, and $\mathbb{D}_{\mathrm{KL}}$ is KL-divergence. The $x$ denotes input prompts drawn from the preference data $\mathcal{D}$. The $j$ indicates the text generated by the reward model, which includes the reasoning trace and final judgement $\hat{l}$. In practice, we use Group Relative Policy Optimization (GRPO) [28] to optimize the objective in Equation 7, the details of which can be find in Appendix C.
@@ 2.3.1 Chain-of-Rubrics (CoR) Rollout
To facilitate the distilled models to proactively generate effective reasoning traces, we design a system prompt as shown in Figure 3 during rollout. Intuitively, reward modeling for general domain (e.g., chat, safety, etc.) and reasoning domain (e.g., math, code, etc.) should focus on different angles. For example, for the chat domain, we may care more about some aspects that can be expressed in textual rubrics (e.g., be polite), yet for the reasoning domain, we usually care more about logical coherence and answer correctness. Based on this intuition, we instruct ${r_{\theta}}$ to classify each preference data sample $\{(x,y_{c},y_{r})\}$ into one of the two <type>: Chat or Reasoning. For each <type>, we prompt ${r_{\theta}}$ to carry out the behavior corresponding to that type step by step: For reasoning tasks, we ask ${r_{\theta}}$ to solve $x$ on its own. During the <eval> phase, $r_{\theta}$ compares $y_{c}$ and $y_{r}$ conditioned on its own </solution> and selects an <answer>. Regarding the Chat type, we instead ask ${r_{\theta}}$ to think about and justify the <rubric> for grading the chat quality (including safety).
@@ 3.2 Main Results
State-of-the-Art Performance. On average, our RM-R1-DeepSeek-Distilled-Qwen-14B model surpasses all previous leading Reward Models (RMs), including INF-ORM-Llama3.1-70B, Nemotron-4-340B-Reward, and GPT-4o, while operating at a much smaller scale. Our 32B models, RM-R1-Qwen-Instruct-32B and RM-R1-DeepSeek-Distilled-Qwen-32B, further extend this lead by a notable margin. The success of RM-R1 is attributable to both our meticulously designed training methodology and the effective scaling of our models, as extensively analyzed in Section 4.1 and Section 4.2. In particular, RM-R1 outperforms existing top-tier ScalarRMs. This highlights the considerable potential of ReasRMs, a category where prior GenRMs have exhibited suboptimal performance and are generally not comparable to their scalar counterparts. In contrast to our structured rollout and distillation with RLVR training strategy, prior critique-based methods have relied heavily on rejection sampling and unstructured, self-generated chain-of-thought (CoT) reasoning from instruct models [22, 35], limiting their reasoning capabilities and leading to inferior performance compared to ScalarRMs. Simultaneously, our comprehensive evaluation indicates that the top-performing scalar models on RewardBench do not consistently achieve state-of-the-art (SOTA) performance; in fact, larger models frequently underperform smaller ones. This evaluation underscores the need for a more comprehensive and systematic approach to RM assessment.
@@ 4.2 Scaling Effects
We then investigate how model performance varies with scale, considering both model size and inference-time compute. In some cases – such as ScalarRMs from InternLM2 [6] and Skywork [20] – the smaller models (7B/8B) outperforms the larger ones (20B/27B), showing no advantage of scaling. In this subsection, we show that this trend does not hold for RM-R1, where scaling brings clear and substantial improvements.
@@ 4.2.2 Inference-time Computation
To ensure a fair comparison, we match the training rollout budget to the inference budget in each setting (i.e., we allow a maximum of $k$ tokens during training for a compute budget of $k$ at inference). All models are trained using GRPO with identical datasets and hyperparameter configurations. Figure 4(b) shows the relationship between compute budget and performance. We observe a clear improvement trend as the inference budget increases. This highlights the benefits of long reasoning chains in reward modeling.
@@ 4.3 Effectiveness of Reasoning Training
Instruct + Distilled + SFT. This approach applies SFT (with respect to the answer directly) after the distillation stage, serving as a direct comparison point with RM-R1 trained with RL.
Instruct + RM-R1 (Distilled + RL). This is the full approach proposed in this paper, following the procedure detailed in Section 2.
@@ 4.4 Case Study
Table 4: Case Study of RM-R1(succinct version). Complete Chatbots and model responses are shown in Table 910, and 11. Cold-Start RL evaluation tends to focus on superficial features (e.g., number of suggestions, relevance), whereas RM-R1 produces high-quality rubrics (e.g., accuracy for factual questions), faithfully follows them, and evaluates responses based on their actual content.
@@ 6 Conclusion and Future Work
This research is based upon work supported DARPA ITM Program No. FA8650-23-C-7316, and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ D.3 Baselines
ScalarRMs. ScalarRMs produce a score for model response directly, predicting preference through single numeric scores without explicit reasoning traces. This category includes models such as Eurus-RM [39], Internlm2 [6] SteerLM-RM [36], Nemotron-RM [2], Tulu-v2.5 [15], Starling-RM [44], ArmoRM [34], Skywork-RM [20], etc. While these models often achieve strong results on well-defined benchmarks, they generally lack interpretability and struggle to capture fine-grained reasoning.
GenRMs. Generative reward models (GenRMs) offer more expressive feedback by producing free-form textual judgments, typically without further training. This includes the widely used LLM-as-a-Judge setup [41], where pretrained language models are prompted to explain and evaluate responses. We also categorize under GenRMs models that directly generate output answers without intermediate reasoning steps. Representative examples include LLaMA [11], Qwen [37], Claude [3], GPT-4o [1, 14], Gemini 1.5 Pro [25], and Skywork-Critic [30]. By leveraging LLMs’ generative capabilities, these models enhance interpretability through natural language rationales and explanations.
ReasRMs. Reasoning-enhanced reward models (ReasRMs) explicitly incorporate reasoning processes before their final judgments, often trained through critiques or chain-of-thought strategies. Notable examples are JudgeLRM [7], Critique-RM [38], DeepSeek-GRM [22], Self-taught Evaluators [35] and our proposed RM-R1 models. These models excel in tasks demanding rigorous reasoning, safety evaluations, and nuanced preference judgments due to their grounding in structured critical thinking.